# Spatio-Temporal Adaptive Diffusion Models for EEG Super-Resolution in Epilepsy Diagnosis

### v3

@@ Abstract
Electroencephalogram (EEG) technology, particularly high-density EEG (HD EEG) devices, are widely used in fields such as neuroscience. HD EEG devices improve the spatial resolution of EEG by placing more electrodes on the scalp, which meet the requirements of clinical diagnostic applications such as epilepsy focus localization. However, this technique faces challenges, such as high acquisition costs and limited usage scenarios. In this paper, spatio-temporal adaptive diffusion models (STADMs) are proposed to pioneer the use of diffusion models for achieving spatial SR reconstruction from low-resolution (LR, 64 channels or fewer) EEG to high-resolution (HR, 256 channels) EEG. Specifically, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which then used as conditional inputs to direct the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance to generate subject-adaptive SR EEG. Experimental results demonstrate that the STADMs significantly enhances the spatial resolution of LR EEG and quantitatively outperforms existing methods. Furthermore, STADMs demonstrate their value by applying synthetic SR EEG to classification and source localization tasks for epilepsy patients, indicating their potential to Substantially boost the spatial resolution of EEG.
{IEEEkeywords}
@@ 1 Introduction
Electroencephalogram (EEG) is a economical, non-invasive neuroimaging technology with high temporal resolution, playing a vital role in cognitive neuroscience research and clinical diagnosis. Unlike other neuroimaging techniques such as positron emission tomography (PET) and functional magnetic resonance imaging (fMRI), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability substantially supports the study of dynamic cognitive processes [1]. Recently, the development of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as sleep monitoring and brain-computer interface (BCI) control [2, 3, 4]. However, these devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [5]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [6, 7, 8]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [9]. These factors collectively lead to difficulties in acquisition and limited application scenarios for HD EEG devices.
@@
The aforementioned observation highlights a significant dilemma in the current clinical diagnosis and treatment using EEG: due to high costs, discomfort from wearing, and other factors, high-resolution (HR) EEG is scarce. Consequently, low-resolution (LR) EEG is insufficient for clinical needs, particularly in the pre-surgery assessment of patients with refractory epilepsy [10]. In such cases, the EEG spatial super-resolution (SR) method offers an effective remedy. This technique enables the reconstruction of HR EEG from LR EEG collected from subjects, thereby enhancing the spatial resolution of the existing EEG acquisition systems. For instance, EEG SR can reconstruct 256-channel EEG from 32-channel counterparts. Notably, EEG SR methods hold substantial potential for clinical applications, providing a cost-effective and efficient method to acquire HR EEG data.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [11]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances [12] and multi-quadratic interpolation methods [13]. However, these methods only utilize the spatial correlations among adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, resulting in suboptimal reconstruction performance for non-stationary EEG signals.
@@
Lately, deep learning approaches have become mainstream for analyzing and enhancing medical images [14, 15, 16, 17]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [18], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [19, 20] to correlate the structural and functional connections of participants’ brains [21]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [22, 23, 24], such as Autoencoders (AE) and Generative Adversarial Networks (GANs) [25, 26, 27]. Existing researches have utilized approaches such as AE and GANs to enhance the spatial resolution of EEG [28, 29]. Although current EEG SR methods based on deep learning have made considerable advancements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [30] have exhibited remarkable generative capabilities in computer vision, leading to a significant increase in the Artificial Intelligence Generated Content (AIGC) [31, 32, 33]. Moreover, diffusion models have demonstrated superior image generation quality compared to GANs, while also providing advantageous features such as comprehensive distribution coverage, extensibility, and stable training targets [34]. Besides, Latent Diffusion Models (LDMs) [35] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Furthermore, SynDiff [36] , Brain Diffuser [37] and DiffMDD [38] have illustrated that diffusion models are applicable in generating medical images, thereby advancing the accuracy of disease diagnosis. Despite this, Applying diffusion models to enhance EEG spatial resolution has not been thoroughly explored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
Figure 1: The architecture of STADMs aimed at generating SR EEG from LR EEG.
@@
In this paper, a novel EEG SR method, called spatio-temporal adaptive diffusion models (STADMs), is proposed to achieve spatial SR reconstruction from LR EEG to HR EEG. Employing a diffusion learning strategy, This approach improves EEG spatial resolution by learning the latent mapping relationship between LR EEG and HR EEG. The primary contributions of this work are summarized as:
@@
    A spatio-temporal adaptive framework based on diffusion models is proposed for EEG SR reconstruction. According to our knowledge, this marks the first time diffusion models are employed to achieve spatial SR reconstruction from LR EEG to HR EEG. This provides an efficient and cost-effective method for obtaining HR EEG, enhancing the portability and usability of HR EEG, and further advancing fields such as epileptic focus localization and BCI technology.
@@
    The spatio-temporal condition module (STCM) is designed to capture the spatio-temporal features of LR EEG. It further uses the generated spatio-temporal conditions to constrain the reverse denoising process, guiding the model to generate SR EEG that conforms to the characteristics of the subjects.
@@
    The multi-scale Transformer denoising module(MTDM) is constructed to map LR EEG to HR EEG. It extracts temporal features at various scales and uses cross-attention-based diffusion Transformer blocks to adaptively modulate the denoising process based on spatio-temporal conditions, addressing the channel-level disparity between LR EEG and HR EEG.
2 Method
2.1 Preliminaries and Problem Statement
@@
2.2 Basic ideas
Learning the mapping relationship from LR EEG to HR EEG is challenging. To ensure that the generated results conform to the subject’s characteristics, the proposed STADMs use the spatio-temporal features of LR EEG as conditions and employs a cross-attention-based conditional injection mechanism to guide the denoising model for progressive denoising restoration. This mechanism embeds the spatio-temporal information of LR EEG into the high-dimensional latent space, indirectly guiding the model to comprehend the potential correlation between LR EEG and HR EEG. This ensures that the synthetic SR EEG is highly consistent with HR EEG in both temporal and spatial domains, addressing the issue of the large channel-level disparity between LR EEG and HR EEG.
As Fig. 1 shows, the entire model framework consists of EEG pre-trained autoencoder, STCM and MTDM. During the training phase of STADMs, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors $z_{0}$. The forward diffusion process progressively derives the state $z_{t}$ of each time step $t$. In the reverse denoising process, the STCM first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTDM takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. To train the MTDM, the Mean Squared Error (MSE) loss is minimized between the predicted and actual noise at each step, ultimately generating SR EEG through progressive denoising. During the sampling phase of STADMs, the deterministic MTDM precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
2.3 Architectures
2.3.1 Pre-trained EEG Autoencoders
@@
2.3.2 Spatio-temporal Condition Module
@@
In EEG signals, each channel represents the signals collected by electrodes placed at specific locations on the brain, recording electrical activity from different brain regions [41, 42]. The spatio-temporal condition module $\tau_{\theta}$ based on Transformer is constructed to capture the temporal correlations between time points and the spatial correlations between EEG channels. This module takes the time series from LR EEG and the spatial positions of the channels as inputs and outputs corresponding encoding vectors $c=\tau_{\theta}(x)$ as conditional information for the reverse denoising process, where $x\in\mathbb{R}^{C\times T}$ represents the LR EEG. By extracting the spatio-temporal features of LR EEG, this module provides richer and more precise contextual information for the denoising model.
@@
As the yellow part of Fig. 1 shows, the STCM is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and process them with the 1D convolution block for the initial extraction of temporal features [43]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition $c$. The 1D convolution block consists of a 1D convolution layer, batch normalization layer and a ReLU activation function. In the Transformer block, a multi-head self-attention (MSA) layer is followed by feed-forward networks, both of which are encapsulated with residual connections and layer normalization.
2.3.3 Multi-scale Transformer Denoising Module

### v4

# Generative AI Enables EEG Super-Resolution via Spatio-Temporal Adaptive Diffusion Learning

@@ Abstract
Electroencephalogram (EEG) technology, particularly high-density EEG (HD EEG) devices, are widely used in fields such as neuroscience. HD EEG devices improve the spatial resolution of EEG by placing more electrodes on the scalp, which meet the requirements of clinical diagnostic applications such as epilepsy focus localization. However, this technique faces challenges, such as high acquisition costs and limited usage scenarios. In this paper, spatio-temporal adaptive diffusion models (STAD) are proposed to pioneer the use of diffusion models for achieving spatial SR reconstruction from low-resolution (LR, 64 channels or fewer) EEG to high-resolution (HR, 256 channels) EEG. Specifically, a spatio-temporal condition module is designed to extract the spatio-temporal features of LR EEG, which then used as conditional inputs to direct the reverse denoising process. Additionally, a multi-scale Transformer denoising module is constructed to leverage multi-scale convolution blocks and cross-attention-based diffusion Transformer blocks for conditional guidance to generate subject-adaptive SR EEG. Experimental results demonstrate that the STAD significantly enhances the spatial resolution of LR EEG and quantitatively outperforms existing methods. Furthermore, STAD demonstrate their value by applying synthetic SR EEG to classification and source localization tasks, indicating their potential to Substantially boost the spatial resolution of EEG.
Index Terms:
@@ I Introduction
Electroencephalogram (EEG) is a economical, non-invasive neuroimaging technology with high temporal resolution, playing a vital role in cognitive neuroscience research and clinical diagnosis. Unlike other neuroimaging techniques such as positron emission tomography (PET) and functional magnetic resonance imaging (fMRI), EEG offers superior temporal resolution and can continuously monitor brain activity in natural environments. This capability substantially supports the study of dynamic cognitive processes [1]. Recently, the development of portable and wireless EEG devices has further improved the flexibility and usability of EEG, enabling its widespread application in emerging neural engineering scenarios such as epileptic seizure detection and brain-computer interface (BCI) control [2, 3, 4]. Besides, these consumer devices can realize real-time remote monitoring of patients through the Internet of Medical Things (IoMT), enhance the early diagnosis and treatment, and improve the performance of real-time applications in intelligent consumer healthcare systems [5]. However, the devices typically have fewer than 64 channels, resulting in lower spatial resolution of the acquired EEG data, which severely limits its potential in clinical diagnosis and brain function imaging research [6]. Currently, high-density (HD) EEG devices improve the spatial resolution of EEG by employing more electrode quantities, thus assisting doctors in performing evaluations or diagnoses in different clinical scenarios. For example, recent studies have shown that non-invasive HD EEG devices can objectively, non-invasively, and accurately determine the epileptic focus area [7, 8, 9]. However, HD EEG devices greatly increase hardware and acquisition costs, their bulkiness and complex wiring prevent subjects from moving their heads freely, restricting their use to controlled environments such as laboratories or hospitals. In addition, due to poor wearing comfort, subjects cannot be monitored for extended periods [10]. Fig. 1 summarizes the respective advantages and limitations of different density EEG devices and reveals difficulties in acquisition and limited application scenarios for HD EEG devices.
Figure 1: Comparative illustration of low-density and high-density EEG Devices: advantages and disadvantages. The picture on the right shows the buddhist monk Barry Kerzin meditating with a high-density EEG device, picture from [11].
@@
The aforementioned observation highlights a significant dilemma in the current clinical diagnosis and treatment using EEG: due to high costs, discomfort from wearing, and other factors, high-resolution (HR) EEG is scarce. Consequently, low-resolution (LR) EEG is inadequate for meeting the demands of emerging neurotechnological applications and clinical requirements, such as the pre-surgery assessment of patients [12]. In such cases, the EEG spatial super-resolution (SR) method offers an effective remedy. Generative AI is able to enhance the spatial resolution of the existing low-density EEG devices by transforming LR EEG into HR EEG, enabling improved brain analysis for applications like BCI, source localization, and sleep monitoring. For instance, EEG SR can reconstruct 256-channel EEG from 32-channel counterparts. Notably, EEG SR methods hold substantial potential for clinical applications, providing a cost-effective and efficient method to acquire HR EEG data.
@@
Currently, in the field of enhancing spatial resolution of EEG, the methodologies predominantly fall into two main categories: channel interpolation based on mathematical models and EEG SR reconstruction utilizing deep learning methods. Channel interpolation methods principally rely on predefined scalp surface models to reconstruct missing channel data using signals from other channels, thereby improving spatial resolution. These methods involve calculating the second-order derivatives of the interpolation function in space to explore the distribution across the electrode channels [13]. Subsequent studies have attempted more precise measurements of the spatial distances between EEG electrodes and more accurate predictions for missing channel data, for example, using inverse distance weighting based on Euclidean distances and multi-quadratic interpolation methods. However, these methods only utilize the spatial correlations among adjacent electrodes and typically struggle to effectively capture and reconstruct the high-frequency and transient components of brain electrical activity, resulting in suboptimal reconstruction performance for non-stationary EEG signals.
@@
Lately, deep learning has become the mainstream method in multiple fields, including the field of Internet healthcare [14, 15, 16] and medical image processing [17, 18]. Specifically, deep learning-based EEG SR technology involves training deep learning models to learn the implicit mappings between LR EEG and HR EEG, and using these trained models to reconstruct SR EEG. Notably, research has introduced EEG SR techniques based on convolution neural networks (CNN) [19], followed by studies proposing deep EEG SR frameworks using graph convolution networks (GCNs) [20] to correlate the structural and functional connections of participants’ brains [21]. Additionally, generative models offer a new learning paradigm for medical image reconstruction [22]. Existing researches have utilized approaches such as Autoencoders (AE) and Generative Adversarial Networks (GANs) [23] to enhance the spatial resolution of EEG. Although current EEG SR methods based on deep learning have made considerable advancements in enhancing EEG spatial resolution, they are still limited in effectively capturing both the temporal dynamics and spatial features of EEG, resulting in limited generalizability and practicality of the models. Moreover, the enhanced spatial resolution achieved by these methods is constrained, with a maximum capability of reconstructing EEG data up to 64-channel level, which may not meet the requirements of specific clinical scenarios.
@@
In clinical diagnostic and therapeutic settings, achieving high-fidelity reconstruction of HR (256 channels) ground truth EEG presents a substantial challenge. The key is accurately depicting the dependencies between LR EEG and HR EEG and addressing the significant channel-level disparity between them. Recently, a novel generative model—diffusion models [24] have exhibited remarkable generative capabilities in computer vision, leading to a significant increase in the Artificial Intelligence Generated Content (AIGC) [25, 26]. Moreover, diffusion models have demonstrated superior image generation quality compared to GANs, while also providing advantageous features such as comprehensive distribution coverage, extensibility, and stable training targets [27]. Besides, Latent Diffusion Models (LDMs) [28] extend diffusion models into latent spaces, utilizing cross-attention mechanisms to achieve refined control in conditional generation tasks. Furthermore, SynDiff [29] and DiffMDD [30] have illustrated that diffusion models are applicable in generating medical images, thereby advancing the accuracy of disease diagnosis. Despite this, Applying diffusion models to enhance EEG spatial resolution has not been thoroughly explored. Therefore, this paper aims to investigate the potential of diffusion models in achieving EEG SR, particularly in enhancing spatial resolution. Through this novel generative model, we hope to provide a novel and effective method for processing and analyzing EEG data, supporting more accurate EEG studies and applications across various neuroscience contexts.
Figure 2: The architecture of STAD aimed at generating SR EEG from LR EEG.
@@
In this paper, a novel EEG SR method, called spatio-temporal adaptive diffusion models (STAD), is proposed to achieve spatial SR reconstruction from LR EEG to HR EEG. Employing a diffusion learning strategy, This approach improves EEG spatial resolution by learning the latent mapping relationship between LR EEG and HR EEG. The primary contributions of this work are summarized as:
@@
    A spatio-temporal adaptive framework based on diffusion models is proposed for EEG SR reconstruction. According to our knowledge, this marks the first time diffusion models are employed to achieve spatial SR reconstruction from LR EEG to HR EEG. This provides an efficient and cost-effective method for obtaining HR EEG, enhancing the portability and usability of HR EEG, and further advancing fields such as source localization and BCI technology.
@@
    The spatio-temporal condition module (STC) is designed to capture the spatio-temporal features of LR EEG. It further uses the generated spatio-temporal conditions to constrain the reverse denoising process, guiding the model to generate SR EEG that conforms to the characteristics of the subjects.
@@
    The multi-scale Transformer denoising module (MTD) is constructed to map LR EEG to HR EEG. It extracts temporal features at various scales and uses cross-attention-based diffusion Transformer blocks to adaptively modulate the denoising process based on spatio-temporal conditions, addressing the channel-level disparity between LR EEG and HR EEG.
II Related Works
In recent years, with the wide application of EEG in neuroscience and clinical diagnosis and treatment, how to improve the spatial resolution of EEG data has become a key problem to be solved. People try to adopt computer technology to enhance the spatial resolution of EEG, including channel interpolation methods based on mathematical models and spatial SR reconstruction methods utilizing deep learning.
II-A Channel Interpolation Methods
The channel interpolation methods mainly utilize the spatial relationship between adjacent electrodes and signal characteristics for the reconstruction of missing electrode data to improve the spatial resolution. To estimate the values of missing channels based on the spatial relationships among electrode locations, Courellis et al.[13] proposed an ellipsoidal geodesic length-based method to reconstruct missing or poor quality EEG signals by synthesizing the signals from neighboring electrodes. Petrichella et al.[31] nvestigated two different interpolation methods to correctly reconstruct artifact-affected channel data and used to map global brain responses after stimulation of specific brain regions of interest. Nouira et al.[32] proposed a multiple quadratic interpolation technique for estimating scalp potential activity in 3-D EEG mapping. In this approach, by replacing the Euclidean distance between two electrodes with the arc length, multiple quadratic interpolation has been shown to be effective in reconstructing SR EEGs for subjects exhibiting various behavioral states.
II-B EEG SR reconstruction utilizing deep learning methods
The deep learning-based spatial super-resolution reconstruction methods aim to learn the mapping relationship between LR EEG and HR EEG using deep learning models. For example, Han et al.[33] employed a deep CNN (DCNN) to enhance the spatial resolution of EEG data and showed that the SR method not only enhances the spatial resolution of EEG, but also improves the signal quality. Tang et al.[21] proposed a deep EEG SR framework termed Deep-EEGSR, which combines a compact convolutional network and a fully-connected filter generation network (FGN). It leverages graph convolutions to adapt to the structural connectivity among EEG channels and employs sample-specific dynamic convolutions to adjust filter parameters to match individual functional connectivity patterns. Saba-Sadiya et al.[34] proposed a deep encoder-decoder network-based approach to automatically interpolate data loss caused by electrode “burst” artifacts in EEG, and used migration learning to enhance the robustness of the model. El-Fiqi et al.[35] developed a weighted gating layer autoencoder for reconstructing lost EEG data. This new variant of the autoencoder focuses on learning the relationships between the variables, which can reveal the interactions between different channels. Corley et al.[36] further explored the application of GANs for EEG spatial super-resolution, aiming to reduce the cost of EEG devices by improving their spatial resolution.
II-C Diffusion Models
In recent years, diffusion models have demonstrated remarkable generative capabilities across various domains, including image, audio, text, and medical data generation. Ho et al.[24] proposed the Denoising Diffusion Probabilistic Model (DDPM), which defines the fundamental components of diffusion models: the forward process and the reverse process, and demonstrated its superiority in image generation. Song et al.[37] introduced the Denoising Diffusion Implicit Model (DDIM), replacing the Markov forward process with a non-Markovian process, thereby improving the sampling speed of diffusion models without compromising the generative performance. Rombach et al.[28] applied diffusion models in the latent space, termed Latent Diffusion Models (LDMs), and utilized cross-attention mechanisms to incorporate conditional information into the underlying UNet architecture, guiding the model’s conditional generation. Gao et al.[38] proposed an implicit diffusion model (IDM) for continuous image super-resolution, which dynamically adjusts the ratio of real information in the low-resolution features to the fine details generated by the diffusion process through a scale-adaptive conditioning mechanism. Peebles et al.[39] proposed a Transformer-based diffusion model framework (DiT), replacing the UNet backbone in LDMs with a Transformer architecture and employing adaptive layer normalization for learning. Ozbey et al.[29] introduced an adversarial diffusion model-based method (SynDiff) for cross-modality translation from Magnetic Resonance Imaging (MRI) to Computed tomography (CT), designing a cycle-consistent architecture for bidirectional modal conversion. Wang et al.[30] proposed a diffusion-based deep learning framework (DiffMDD) for Major Depressive Disorder (MDD) diagnosis using EEG, incorporating a forward diffusion noisy training module to extract noise-irrelevant features for improved robustness, and a reverse diffusion data augmentation module to increase data size and diversity for learning generalized features.
III Method
III-A Preliminaries and Problem Statement
@@
III-B Basic ideas
Learning the mapping relationship from LR EEG to HR EEG is challenging. To ensure that the generated results conform to the subject’s characteristics, the proposed STAD use the spatio-temporal features of LR EEG as conditions and employs a cross-attention-based conditional injection mechanism to guide the denoising model for progressive denoising restoration. This mechanism embeds the spatio-temporal information of LR EEG into the high-dimensional latent space, indirectly guiding the model to comprehend the potential correlation between LR EEG and HR EEG. This ensures that the synthetic SR EEG is highly consistent with HR EEG in both temporal and spatial domains, addressing the issue of the large channel-level disparity between LR EEG and HR EEG.
As Fig. 2 shows, the entire model framework consists of EEG pre-trained autoencoder, STC and MTD. During the training phase of STAD, the HR EEG ground truth is first input into the EEG pre-trained encoder to obtain the corresponding latent vectors $z_{0}$. The forward diffusion process progressively derives the state $z_{t}$ of each time step $t$. In the reverse denoising process, the STC first encodes the time series and channel spatial distribution matrix from LR EEG into the spatio-temporal representation conforming to the subject, and concatenates it with time embedding as the conditioning vector. MTD takes the conditioning vectors and noise sampled from the standard normal distribution as input, implements the conditional guidance mechanism through the cross-attention layer, and predicts the noise at each step of the reverse process. To train the MTD, the Mean Squared Error (MSE) loss is minimized between the predicted and actual noise at each step, ultimately generating SR EEG through progressive denoising. During the sampling phase of STAD, the deterministic MTD precisely generates SR EEG matching the input LR EEG by progressively sampling using the spatio-temporal conditional information of LR EEG and noise sampled from the Gaussian distribution.
III-C Architectures
III-C1 Pre-trained EEG Autoencoders
@@
III-C2 Spatio-temporal Condition Module
@@
In EEG signals, each channel represents the signals collected by electrodes placed at specific locations on the brain, recording electrical activity from different brain regions [42]. The spatio-temporal condition module $\tau_{\theta}$ based on Transformer is constructed to capture the temporal correlations between time points and the spatial correlations between EEG channels. This module takes the time series from LR EEG and the spatial positions of the channels as inputs and outputs corresponding encoding vectors $c=\tau_{\theta}(x)$ as conditional information for the reverse denoising process, where $x\in\mathbb{R}^{C\times T}$ represents the LR EEG. By extracting the spatio-temporal features of LR EEG, this module provides richer and more precise contextual information for the denoising model.
@@
As the yellow part of Fig. 2 shows, the STC is mainly composed of a spatial position embedding layer, a 1D convolution block, and a Transformer block. More specifically, the channel spatial embedding layer encodes the channel spatial coordinates of LR EEG, outputting a set of spatial features. Then, we partition the input time series into different patches and process them with the 1D convolution block for the initial extraction of temporal features [43]. Finally, the Transformer block concatenates the temporal features with position encoding and channel spatial embeddings to output the final spatio-temporal condition $c$. The 1D convolution block consists of a 1D convolution layer, batch normalization layer and a ReLU activation function. In the Transformer block, a multi-head self-attention (MSA) layer is followed by feed-forward networks, both of which are encapsulated with residual connections and layer normalization.
III-C3 Multi-scale Transformer Denoising Module